OpenStudy (anonymous):

Your gross pay is $1,843.45. Your involuntary deductions are FICA (7.65%), federal withholding (9%), and state withholding (6.5%). How much are you allowed for housing and fixed expenses? (1 point)$516.17 $663.64 $510.01 $396.68

OpenStudy (perl):

Did you try to change percents to decimal equivalents, and multiply. 1843.45 - .0765 *1843.45 - .09*1843.45 - .065 *1843.45

OpenStudy (perl):

None of the answers match that, so I am probably doing something wrong. Do you have a formula from a book to use?

OpenStudy (anonymous):

No, the only slight thing they gave us is to never go over 36% on housing/expenses

OpenStudy (perl):

ok so the net income (the money you can take home after taxes etc) is 1843.45 - .0765 *1843.45 - .09*1843.45 - .065 *1843.45 =1416.691325 now take 36% of that

OpenStudy (perl):

ok now it matches one of the solution

OpenStudy (anonymous):

Alright.. So 510.01?

OpenStudy (anonymous):

Could you help me with another problem or so? I can open it in a different thing.

OpenStudy (perl):

we can do it here

OpenStudy (perl):

was 510 correct?

OpenStudy (anonymous):

I won't know until I submit it, and I'm not done yet.

OpenStudy (anonymous):

My apologies.

OpenStudy (perl):

we need the definition of debt ratio

OpenStudy (perl):

what book are you studying from

OpenStudy (anonymous):

Uh, I don't know. I'm online schooled.

OpenStudy (perl):

do you have any notes on 'debt ratio'

OpenStudy (anonymous):

One moment, let me look. Sorry.

OpenStudy (anonymous):

"If you have five credit cards, each of them with a $1,000.00 credit limit, you have $5,000.00 of available credit. Now assume you have charged a total of $5,000.00 on them. That means you have a 1:1 ratio, or 100% debt. Big problem! That tells the credit reporting agencies that you're completely maxed out, and it's unlikely a bank or credit card company will extend you any loans until you get your finances under control."

OpenStudy (anonymous):

So far this is all I've seen on it, throughout the entire unit.

OpenStudy (perl):

so the ratio we are using is the balance-to-limit ratio that means if you have 100 dollar balance and a 1000 dollar credit card limit you have 100/1000 = .10 or 10% ratio. In our problem the total debt is 1800 + 2600 + 3500 = 7900 the limit is also 7900. so we start with balance to limit of 100% paying off and closing any of the three credit cards will still leave the other two credit balance at maximum , so the debt remains the same

OpenStudy (perl):

because you will still have 100 % debt on the other two cards

OpenStudy (anonymous):

..Alright, I understand now! Thank you very much.